find the general term of the sequence, starting with n = 1. Determine whether the sequence converges and if so find its limit. If the sequence diverges, indicate that using the checkbox.


3, 3/19, 3/19^2, 3/19^3...



1
Expert's answer
2021-12-01T17:22:41-0500

3, 3/19, 3/19^2, 3/19^3...

Given sequence is a GP whose first term "(a)" is 3 and common ratio "(r)" is 1/19.

"n^{th}" term of a GP is given by: "a_n=a.r^{n-1}"

"a_n=\frac{3}{19^{n-1}}"

"\lim_{n\to\infty} a_n= \lim_{n\to\infty} \frac{3}{19^{n-1}}=0"